OpenStudy (bananas1234):

solve each equation. Write irrational numbers in simplest radical form. Describe the strategy you used to get your solution and tell why you chose that strategy. 1. x^2 + 4 = 0 2. x^2 - 6x + 1 = 0

OpenStudy (anonymous):

Solve for x

OpenStudy (bananas1234):

@iambatman

OpenStudy (jack1):

1. using reverse FOIL: (x+2)(x-2) = x^2 + 2x - 2x + 4 = x^2 + 4 so if x^2 + 4 = 0 (x+2)(x-2) =0 so either x+2 = 0 or x - 2 = 0 if it's x+2 = 0... then x = -2 if it's x-2 = 0... then x = 2 so either x = 2 or x = -2 is this making sense? can you do Q2?

OpenStudy (bananas1234):

i think so, thanks!

OpenStudy (jack1):

sweet ;)